# XAMI - A Benchmark Dataset for Artefact Detection in XMM-Newton Optical Images

### v1

@@ Abstract
The presented method and dataset will advance artefact detection in astronomical observations by providing a reproducible baseline. All code and data are made available publicly111https://github.com/ESA-Datalabs/XAMI-model,222https://github.com/ESA-Datalabs/XAMI-dataset.
@@ 3 Results
Overall | 90.1 | 73.5 | 55.4 |
Small | 82.6 | 59.6 | 48.3 |
Medium | 91.2 | 75.3 | 57.1 |
Large | 84.6 | 63.4 | 58.0 |
CR | 97.6 | 94.9 | 70.2 |
SR | 94.8 | 77.7 | 50.6 |
SL | 91.8 | 75.5 | 57.3 |
ROS | 82.3 | 55.2 | 48.2 |
Other | 83.9 | 64.1 | 50.6 |
@@ 4 Discussion
In our study, we enhanced artefact detection and segmentation in XMM-Newton images by integrating CNNs with ViT-based models, significantly boosting accuracy and reducing false positives in astronomical analysis. We combined traditional YOLO models for bounding box predictions with advanced SAM models for zero-shot segmentation, demonstrating the benefits of diverse neural network strategies in addressing complex image processing challenges. Despite improvements, high variation in exposure times and intensity levels in space imagery necessitate further model refinement tailored for astronomical missions. Additionally, these variations pose challenges in dataset annotation, eventually making it difficult to establish clear thresholds for distinguishing artefacts from the background. The XAMI average end-to-end inference time per image containing annotations is , suitable for medium to large image data (up until hundreds of thousands of images similar to ours) and applications which do not particularly require instant real-time processing. The SAM heavy architecture still represents a bottleneck for prediction, with image.
Acknowledgements. The authors acknowledge the contribution of Inès Perez, Léa Zuili, Simon Astarita to dataset annotations. This publication uses data generated via the Roboflow.com and Zooniverse.org platforms.

### v2

@@ Abstract
[…]
@@ 3 Results
| Overall | 79.1 | 63.6 | 46.3 |
| Small | 68.7 | 50.6 | 39.3 |
| Medium | 80.0 | 63.7 | 45.3 |
| Large | 80.8 | 66.0 | 53.5 |
| CR | 96.8 | 93.4 | 64.7 |
| SR | 87.7 | 70.8 | 46.2 |
| SL | 84.6 | 66.7 | 50.9 |
| ROS | 60.4 | 34.7 | 33.3 |
| Other | 66.0 | 52.6 | 36.1 |
@@ 4 Discussion
In our study, we enhanced artefact detection and segmentation in XMM-Newton images by integrating CNNs with ViT-based models, significantly boosting accuracy and reducing false positives in astronomical analysis. We combined traditional YOLO models for bounding box predictions with advanced SAM models for zero-shot segmentation, demonstrating the benefits of diverse neural network strategies in addressing complex image processing challenges. Despite improvements, high variation in exposure times and intensity levels in space imagery necessitate further model refinement tailored for astronomical missions. Additionally, these variations pose challenges in dataset annotation, eventually making it difficult to establish clear thresholds for distinguishing artefacts from the background. The XAMI average end-to-end inference time per image containing annotations is , suitable for medium to large image data (up until hundreds of thousands of images similar to ours) and applications which do not particularly require instant real-time processing. The SAM heavy architecture still represents a bottleneck for prediction, with image. We plan to expand our dataset with additional observations from various space missions to enhance model performance. Our goal is to benchmark our methods against other techniques in artefact detection and segmentation, while allowing the flexibility to choose or replace detectors and segmentors with novel state-of-the-art models. We will explore practical applications, such as faster data processing for space missions, and integrate our implementations into existing astronomical data processing systems. Feedback from astronomers will be crucial in further refining our approach.
[…]